# Dephasing and error dynamics affecting a singlet-triplet qubit during coherent spin shuttling

### v1

@@ Abstract
Quantum information transport over micron to millimeter scale distances is critical for the operation of practical quantum processors based on spin qubits. One method of achieving a long-range interaction is by coherent electron spin shuttling through an array of silicon quantum dots. In order to execute many shuttling operations with high fidelity, it is essential to understand the dynamics of qubit dephasing and relaxation during the shuttling process in order to mitigate them. However, errors arising after many repeated shuttles are not yet well documented. Here, we probe decay dynamics contributing to dephasing and relaxation of a singlet-triplet qubit during coherent spin shuttling over many $N$ repeated shuttle operations. We find that losses are dominated by magnetic dephasing for small $N<10^{3}$ and by incoherent shuttle errors for large $N>10^{3}$. Additionally, we estimate shuttle error rates below $1\times 10^{-4}$ out to at least $N=10^{3}$, representing an encouraging figure for future implementations of spin shuttling to entangle distant qubits.
@@ I Introduction
Spin qubits are a compelling candidate for the foundation of a quantum register network, having demonstrated single and two-qubit gate fidelities well over 99% Yoneda2018 ; Mills2022 . By reducing the spinful 29Si nuclear background in silicon, isotopic enrichment enables coherent spin qubits to be hosted in foundry-compatible Zwerver2022 ; Neyens2024 gate-defined quantum dots (QDs) vanderWiel2006 ; Burkard2023 , streamlining integration with existing and rapidly advancing industrial silicon technology. Since spin qubit operations are carried out using short-range interactions often involving nearest neighbors Noiri2022 ; Watson2018 ; Xue2022 ; Sigillito2019 , a scalable quantum computing architecture relying on long-range quantum information transfer using spin qubits could present a challenge. One solution is to build a remote qubit registry that facilitates information exchange over broad physical scales, leaving qubit operations to be carried out at dedicated local nodes VanMeter2013 ; Vandersypen2017 .
@@
Linking quantum information between qubit nodes can be achieved in multiple ways. One example is by interfacing distant stationary qubit sites through an intermediate messenger, such as through virtual microwave photons to mediate remote spin-spin interactions in quantum transducers Mi2018 ; Borjans2020 ; Harvey-Collard2022 . Another approach is to physically move the spin qubit; for example, by using surface acoustic waves to form a moving QD Takada2019 ; Jadot2021 , or by shuttling the spin controllably through an array of gate-defined electrostatic potential traps Flentje2017 ; Fujita2017 ; vanRiggelen-Doelman2023 ; Langrock2023 . In the latter, the ability to perform many repeated shuttling operations with a low error rate is key to maintaining long-range qubit entanglement. A firm understanding of dephasing and relaxation sources during shuttling is therefore required. At present, these dynamics are not well quantified, especially out to the large number of shuttles needed to perform long-range interactions. For a spin qubit device composed of millions of qubits, for example, arranged on a crossbar array Borsoi2024 , thousands to tens of thousands of shuttling operations will need to be performed to link sites on either end of the processor.
@@
In this work, we categorize the main error sources and dynamics during electron spin shuttling by observing a singlet-triplet qubit over many repeated coherent shuttles of a single electron spin in a gate-defined Si/SiGe triple QD device. We focus first on expected shuttling-independent loss mechanisms – inhomogeneous dephasing and spin relaxation – and use this baseline to classify the nature of errors dependent on the number of shuttles ($N$). From Ramsey decay measurements carried out on the stationary qubit and the qubit during shuttling, we attribute the dominant error source occurring during shuttling for $N<10^{3}$ to inhomogeneous dephasing arising from residual nuclear spins. We corroborate this finding by using dynamical decoupling to reduce magnetic noise and show it is independent of $N$, while also discovering the presence of nontrivial noise dynamics. From relaxation measurements, we define a bounding timescale expected of incoherent error processes in the qubit. We find, after implementing many repeated shuttling operations, that shuttling errors are incoherent but don’t contribute appreciably to losses until $N>10^{3}$, meaning that errors due to shuttling are small. Understanding the various noise mechanisms encountered during shuttling can aid the development of targeted mitigation strategies that will enable the completion of a large number of shuttles, allowing an opportunity for linking localized Morton2008 ; Pla2013 ; Witzel2022 , and even deterministic Jakob2022 , qubit nodes on a sparse qubit array.
@@ I.1 Device overview and operation
We operate the device with four electrons in the QD chain, where two low-energy electrons under P0 form a filled-shell singlet and do not contribute significantly to the spin dynamics of the system Higginbotham2014 ; Harvey-Collard2017 . Fig. 1c shows an energy diagram of the remaining two-spin system illustrating the relationship between the singlet (|$\ket{S}$) and triplet (|$\ket{T_{0}},\leavevmode\nobreak\ \ket{T_{+}},\leavevmode\nobreak\ \ket{T_{-}}$) states in a small magnetic field. We employ a singlet-triplet ($S$-$T_{0}$) qubit Petta2005 , where the qubit is encoded in the $m=0$ subspace of the 4-level system defined by linear combinations of the $\ket{\uparrow\downarrow}$ and $\ket{\downarrow\uparrow}$ states. Its logical states are defined as the eigenstates of the system in the limit of large exchange:
@@
A schematic of the pulse sequence used to initialize and measure the qubit is depicted in Fig. 1e, where we use an enhanced latching mechanism Harvey-Collard2018 of spin-to-charge conversion through Pauli spin blockade (PSB) to measure the qubit spin state Johnson2005 ; Lai2011 ; Zhao2019 ; Jang2020 . Two-axis control of the qubit is achieved through baseband voltage pulses to modify barrier and QD chemical potentials and toggle between control regimes dominated by the exchange interaction, $J$, and the difference in electron Zeeman energy splitting between electrons in two QDs, $\Delta E_{\mathrm{Z}}$ (Fig. 1d). We use voltage pulses applied to the B1 barrier gate and adjacent P0 and P1 QD gates at the exchange symmetric operating point Martins2016 ; Reed2016 to achieve rotations between $\ket{\uparrow\downarrow}$ and $\ket{\downarrow\uparrow}$. When the electrons are well isolated and exchange is turned off, |$\ket{S}$ to |$\ket{T_{0}}$ rotations are driven by Jock2022
@@
where $g$ is the $g$-factor, $\mu_{\mathrm{B}}$ is the Bohr magneton, and $B$ is the magnetic field. Differences in local magnetic fields or local $g$-factors both contribute to $\Delta E_{\mathrm{Z}}$. The unique $B$ at each QD arises from external magnetic field gradients, or from the presence of residual spinful nuclei forming a random but slowly-varying Overhauser field Koppens2005 ; Coish2008 . The difference in $g$-factor at each QD comes from variations in interface disorder across the Si/SiGe interface giving rise to local spin-orbit interactions Prada2011 ; Jiang2012 .
@@ II.1 Coherent spin transfer
Two established approaches for spin shuttling Langrock2023 are conveyor mode Mills22019 ; Seidler2022 ; Xue2023 ; Struck2024 and bucket brigade mode Flentje2017 ; Fujita2017 ; Noiri2022 ; vanRiggelen-Doelman2023 ; Zwerver2023 . Shuttling in this work is accomplished by a barrier pulse-assisted bucket brigade mode, in which we leverage individual gate voltage control to tailor the electrostatic potential landscape and move the spin between QDs. We prepare a |$\ket{S}$ state and transfer one electron of the entangled spin pair from QD1 to QD2 (charge states (3,1,0) and (3,0,1), respectively) by detuning the voltages $V_{\mathrm{P1}}$ and $V_{\mathrm{P2}}$ (ramping in opposite directions), while simultaneously pulsing $V_{\mathrm{B2}}$ to lower the potential barrier and facilitate adiabatic transfer (Fig. 2a) (see discussion on Supplementary Fig. 1 for more details). We hold $V_{\mathrm{B1}}$ fixed to maintain negligible residual $J$ with the electron in QD0 and ensure rotations about the Bloch sphere are facilitated by the dominant $\Delta E_{\mathrm{Z}}$ interaction.
@@
We first demonstrate spin transfer from one QD to another by examining qubit rotations in the region near the (3,1,0)-(3,0,1) charge transition. We apply an external magnetic field of $B_{0}=50$ mT in the plane of the Si quantum well to intentionally drive strong $S$-$T_{0}$ rotations in both charge regions, showing the spin is coherently transferred. Fig. 2b shows singlet return probability following a 6 $\mu$s wait time at different QD1-QD2 detunings. This plot has four regions of interest describing different charge configurations. The bands in (3,1,0) and (3,0,1) describe differences in acquired singlet phase within each charge configuration due to changes in $g$ modifying $\Delta E_{\mathrm{Z}}$ as QD1 and QD2 are shaped by applied voltages, and the local interface disorder sampled by the electron varies Liu2021 . The charge transition between the (3,1,0) and (3,0,1) appears as a diagonal strip of constant $\sim$50singlet return probability with no acquired phase accumulation visible, indicating a fully dephased qubit. A shift in coherent singlet return on either side of the transition indicates the spin is transferred from one isolated QD to the other (see discussion on Supplementary Fig. 2 for more details). The low singlet return probability and absence of acquired phase bands in (3,1,1) ((3,0,0)) is due to the addition (subtraction) of an electron to the system which rapidly destroys qubit coherence. Altogether, these four delineated regions map out the charge space where spin shuttling operations will be performed and show that the entangled electron spin may be coherently transferred between QD1 and QD2.
@@ II.2 Spin coherence and sources of dephasing
We first consider dephasing in the case of negligible external magnetic field ($B_{0}\approx 0$ mT) such that dephasing is expected to be dominated by fluctuations in the local Overhauser field Petta2005 , where no $S$-$T_{0}$ rotations are observed. We fit the data to a decay profile of the form
@@
where $A$ is the amplitude, $y_{0}$ is the vertical offset, $T_{2}^{*}$ is the inhomogeneous dephasing time, and $\alpha$ is the decay exponent. The resulting dephasing times and fit uncertainties are $T_{2}^{*}(3,1,0)=4.60\pm 0.02\leavevmode\nobreak\ \mu$s, $T_{2}^{*}(3,0,1)=3.86\pm 0.02\leavevmode\nobreak\ \mu$s, and $T_{2}^{*}(\mathrm{transition})=5.12\pm 0.03\leavevmode\nobreak\ \mu$s, and are consistent with previously measured quantities for QDs in 800 ppm 29Si Eng2015 ; Jock2018 ; Neyens2024 . The variation among $T_{2}^{*}$s suggest local sources of dephasing affect the spin in each QD, including the wavefunctions of nearby 29Si and 73Ge. We interpret the longer $T_{2}^{*}(\mathrm{transition})$ as the result of a process like motional narrowing – effectively averaging out magnetic noise acquired from the different nuclear distributions when the spin is hybridized between QD1 and QD2 Langrock2023 . The resulting decay exponent fits, $\alpha(3,1,0)=1.92\pm 0.02$, $\alpha(3,0,1)=2.16\pm 0.02$, and $\alpha(\mathrm{transition})=2.12\pm 0.02$, resemble Gaussian profiles as expected of inhomogeneous nuclear spin dephasing Petta2005 .
Next, we apply a small external field of $B_{0}=50$ mT to amplify and study the effect of spin-orbit-driven charge noise acting on $\Delta E_{\mathrm{Z}}$. Clear $S$-$T_{0}$ oscillations are observed and fit to a decaying sine function of the form
@@
where $f$ is the $S$-$T_{0}$ rotation frequency. The resulting dephasing times are $T_{2}^{*}(3,1,0)=3.82\pm 0.12\leavevmode\nobreak\ \mu$s, $T_{2}^{*}(3,0,1)=4.09\pm 0.03\leavevmode\nobreak\ \mu$s, and $T_{2}^{*}(\mathrm{transition})=1.08\pm 0.01\leavevmode\nobreak\ \mu$s. We attribute the fast decay observed at the charge transition to the hybridization of the electron wavefunction split between QD1 and QD2 Yoneda2021 , which is sensitive to charge noise. Charge noise arises from rapid fluctuations in spin-orbit interactions at the Si/SiGe interface that are intensified in an applied external magnetic field.
Additionally, we observe changes to $\alpha$ that imply changes to the noise spectrum. The resulting decay exponent fits are $\alpha(3,1,0)=0.90\pm 0.06$, $\alpha(3,0,1)=2.12\pm 0.06$, and $\alpha(\mathrm{transition})=1.14\pm 0.02$, ranging from exponential to Gaussian forms. Since it is the noise spectrum and mechanisms that dictate the observed decay profile Stano2022 , the spread in $\alpha$ suggests that different localized noise dynamics, in addition to residual nuclear spins, contribute to dephasing in each QD. One possible source known to exhibit non-Gaussian decay behavior is random telegraph noise originating from charge traps Culcer2009 , where the observed charge noise is often attributed to two-level systems (TLSs) that are distributed nonuniformly across the Si/SiGe interface Connors2019 . This is consistent with our assessment that charge noise dominates the fast $T_{2}^{*}(\mathrm{transition})$, and suggests an increased susceptibility to charge noise in QD1 compared to QD2, since $\alpha(3,1,0)$ is exponential and $\alpha(3,0,1)$ is Gaussian. Overall, we observe qubit dephasing to be primarily driven by fluctuations in the Overhauser fields generated by residual nuclear spins at 0 mT, and a combination of nuclear spins and charge noise at 50 mT.

### v2

@@ Abstract
Quantum information transport over micron to millimeter scale distances is critical for the operation of practical quantum processors based on spin qubits. One method of achieving a long-range interaction is by coherent electron spin shuttling through an array of silicon quantum dots. In order to execute many shuttling operations with high fidelity, it is essential to understand the dynamics of qubit dephasing and relaxation during the shuttling process in order to mitigate them. However, errors arising after many repeated shuttles are not yet well documented. Here, we probe decay dynamics contributing to dephasing and relaxation of a singlet-triplet qubit during coherent spin shuttling over many $N$ repeated shuttle operations. We find that losses are dominated by magnetic dephasing for small $N<10^{3}$ and by incoherent shuttle errors for large $N>10^{3}$. Additionally, we estimate shuttle error rates below $10^{-4}$ out to at least $N=10^{3}$, representing an encouraging figure for future implementations of spin shuttling to entangle distant qubits.
@@ I Introduction
Spin qubits are a compelling candidate for the foundation of a quantum register network, having demonstrated single and two-qubit gate fidelities over 99% Yoneda2018 ; Mills2022 ; Noiri2022 ; Xue2022 . By reducing the spinful 29Si nuclear background in silicon, isotopic enrichment enables coherent spin qubits to be hosted in foundry-compatible gate-defined quantum dots (QDs) Ha2022 ; Neyens2024 ; Huang2024 , streamlining integration with existing and rapidly advancing industrial silicon technology. Since spin qubit operations are carried out using short-range interactions often involving nearest neighbors Veldhorst2015 ; Sigillito2019 , a scalable quantum computing architecture relying on long-range quantum information transfer using spin qubits could present a challenge. One solution is to build a remote qubit registry that facilitates information exchange over processor-length scales by carrying out qubit operations at dedicated local nodes linked by interconnected hardware VanMeter2013 ; Vandersypen2017 .
@@
Linking quantum information between qubit nodes can be achieved in multiple ways. One example is by interfacing distant stationary qubit sites through an intermediate messenger, such as through virtual microwave photons to mediate remote spin-spin interactions in quantum transducers Mi2018 ; Borjans2020 ; Harvey-Collard2022 . Another approach is to physically move the spin qubit, for example, by using surface acoustic waves to form a moving QD Takada2019 ; Jadot2021 , or by shuttling the spin controllably through an array of gate-defined electrostatic potential traps Fujita2017 ; vanRiggelen-Doelman2024 ; Langrock2023 ; DeSmet2024 . In the latter, the ability to perform many repeated shuttling operations with a low error rate is key to maintaining long-range qubit entanglement. A firm understanding of error sources during shuttling is therefore required. For a spin qubit device composed of millions of qubits, for example, arranged on a crossbar array Li2018 ; Borsoi2024 or a sparse qubit node array Vandersypen2017 ; Kunne2024 , thousands to tens of thousands of shuttling operations may need to be performed to link sites on either end of a macroscopic processor. At present, error dynamics are not well quantified out to the large number of shuttles needed to establish these long-range interactions.
@@
In this work, we categorize the main error sources and dynamics during electron spin shuttling by observing a singlet-triplet qubit over many repeated coherent shuttles of a single electron spin in a gate-defined Si/SiGe triple QD device. We focus first on expected shuttling-independent loss mechanisms – inhomogeneous dephasing and spin relaxation – and use this baseline to classify the nature of errors dependent on the number of shuttles ($N$). From Ramsey decay measurements carried out on the stationary qubit and the qubit during shuttling, we attribute the dominant error source occurring during shuttling for $N<10^{3}$ to inhomogeneous dephasing arising from residual nuclear spins. We corroborate this finding by using dynamical decoupling to reduce magnetic noise and show it is independent of $N$, while also discovering the presence of nontrivial noise dynamics. From relaxation measurements, we define a bounding timescale expected of incoherent error processes in the qubit. We find, after implementing many repeated shuttling operations, that additional shuttling errors are incoherent but do not contribute appreciably to losses until $N>10^{3}$, meaning that errors due to shuttling are small.
Figure 1: Singlet-triplet qubit in a triple quantum dot device. a. 3D profile of the device showing the 28Si quantum well (light blue) buried within SiGe (dark blue), overlapping metal gates (maroon) and oxide (gray) (above). 2D overview of the device with relevant gates composing the SET and the shuttling QDs encircled (below). b. Cartoon schematic of the electrostatic potential in the shuttling QDs depicting the charge transition between (4,0,0) and (3,1,0). The electron reservoir is shown under accumulation gate AC0 in purple. Electron spins ordered on energy levels (light blue horizontal lines) are moved between adjacent QDs by applying voltages to the plunger (P) and barrier (B) gates in light blue. Voltages on other gates in dark blue are nominally static. c. Energy diagram of the $S$-$T_{0}$ qubit in an applied magnetic field. $J$ is defined between |$\ket{S}$ and |$\ket{T_{0}}$ states. $\Delta E_{\mathrm{Z}}$ separates the $\ket{\uparrow\downarrow}$ and $\ket{\downarrow\uparrow}$ states. $\ket{\uparrow\downarrow}$ is arbitrary in permutation and cannot be distinguished from $\ket{\downarrow\uparrow}$, but is used throughout for simplicity. d. $S$-$T_{0}$ Bloch sphere corresponding to the energy diagram in c. showing orthogonal energy components $J$ and $\Delta E_{\mathrm{Z}}$ e. Charge stability map formed by varying P0 and P1 gate voltages showing charge transitions between QD0 and QD1. The qubit is reset in (3,0,0) for 140 $\mu$s (R), then loaded into (4,0,0) (L), rolled near the (4,0,0)–(3,1,0) PSB region (approximate location outlined in pink) to prepare for transfer (P), and then transferred by rapid adiabatic passage to (3,1,0) (Q). The spin state is measured by latched PSB at the (4,0,0)-(4,1,0) boundary (M).
[…]
@@ I.1 Device overview and operation
We operate the device with four electrons in the QD chain, where two low-energy electrons under P0 form a filled-shell singlet and do not contribute significantly to the spin dynamics of the system. Fig. 1c shows an energy diagram of the remaining two-spin system illustrating the relationship between the singlet (|$\ket{S}$) and triplet (|$\ket{T_{0}},\leavevmode\nobreak\ \ket{T_{+}},\leavevmode\nobreak\ \ket{T_{-}}$) states in a small magnetic field. We employ a singlet-triplet ($S$-$T_{0}$) qubit Petta2005 , where the qubit is encoded in the $m=0$ subspace of the 4-level system defined by linear combinations of the $\ket{\uparrow\downarrow}$ and $\ket{\downarrow\uparrow}$ states. Its logical states are defined as the eigenstates of the system in the limit of large exchange:
@@
A schematic of the pulse sequence used to initialize and measure the qubit is depicted in Fig. 1e, where we use an enhanced latching mechanism of spin-to-charge conversion through Pauli spin blockade (PSB) to measure the qubit spin state Harvey-Collard2018 . Two-axis control of the qubit is achieved through baseband voltage pulses to modify barrier and QD chemical potentials and toggle between control regimes dominated by the exchange interaction, $J$, and the difference in electron Zeeman energy splitting between electrons in two QDs, $\Delta E_{\mathrm{Z}}$ (Fig. 1d). We use voltage pulses applied to the B1 barrier gate and adjacent P0 and P1 QD gates at the exchange symmetric operating point Martins2016 ; Reed2016 to achieve rotations between $\ket{\uparrow\downarrow}$ and $\ket{\downarrow\uparrow}$. When the electrons are well isolated and exchange is turned off, |$\ket{S}$ to |$\ket{T_{0}}$ rotations are driven by Jock2018
@@
where $g$ is the $g$-factor, $\mu_{\mathrm{B}}$ is the Bohr magneton, and $B$ is the magnetic field. Differences in local magnetic fields or local $g$-factors both contribute to $\Delta E_{\mathrm{Z}}$. The unique $B$ at each QD arises from external magnetic field gradients, or from the presence of residual spinful nuclei forming a random but slowly-varying Overhauser field Koppens2005 . The difference in $g$-factor at each QD comes from variations in interface disorder across the Si/SiGe interface giving rise to local spin-orbit interactions Jock2018 .
@@ II.1 Coherent spin transfer
Two established approaches for spin shuttling are conveyor mode Seidler2022 ; Xue2024 ; Struck2024 and bucket brigade mode Flentje2017 ; Fujita2017 ; Mills2019 ; Noiri2022 ; vanRiggelen-Doelman2024 ; Zwerver2023 . Shuttling in this work is accomplished by a barrier pulse-assisted bucket brigade mode, in which we leverage individual gate voltage control to tailor the electrostatic potential landscape and move the spin between QDs. We prepare a |$\ket{S}$ state in QD0 and QD1 and transfer one electron of the entangled spin pair from QD1 to QD2 (charge states (3,1,0) and (3,0,1), respectively) by detuning the voltages $V_{\mathrm{P1}}$ and $V_{\mathrm{P2}}$ (ramping in opposite directions), while simultaneously pulsing $V_{\mathrm{B2}}$ to lower the potential barrier and facilitate adiabatic transfer (Fig. 2a). In principle, this shuttling method is agnostic to spin qubit encoding, provided individual spin control is possible. The tunnel coupling between QD1 and QD2 is estimated around 52 GHz with negligible probability of completing a nonadiabatic transition (see discussion on Supplementary Fig. 2b for more details). We hold $V_{\mathrm{B1}}$ fixed to maintain negligible residual $J$ with the electron in QD0 and ensure rotations about the Bloch sphere are driven by the dominant $\Delta E_{\mathrm{Z}}$ interaction (see Supplementary Table S2 for a voltage level table).
Figure 2: Shuttling and inhomogeneous dephasing. a. Cartoon representation of an electron being shuttled from QD1 to QD2. b. Singlet return probability (|$\ket{S}$%) during shuttling the detuning space surrounding the (3,1,0)-(3,0,1) charge transition. Three colored dots represent P1-P2 detunings corresponding to measurements shown in c. c. Inhomogeneous dephasing of the singlet return probability over time spent at three detunings: (3,1,0) (gold), (3,0,1) (pink) and at the charge transition (violet). Data at $B_{0}\approx 0$ (50) mT are shown as small translucent (large opaque) circles, with decay envelope fits to Eq. 3 overlaid as dashed (solid) lines. Curves are vertically offset for clarity. Full fit parameter values and uncertainties are tabulated in Supplementary Table 3.
@@
We first demonstrate spin transfer from one QD to another by examining qubit rotations in the region near the (3,1,0)-(3,0,1) charge transition. We apply an external magnetic field of $B_{0}=50$ mT in the plane of the Si quantum well to intentionally drive strong $S$-$T_{0}$ rotations in both charge regions, showing the spin is coherently transferred. Fig. 2b shows singlet return probability following a 6 $\mu$s wait time at different QD1-QD2 detunings. This plot has four regions of interest describing different charge configurations. The bands in (3,1,0) and (3,0,1) describe differences in acquired singlet phase within each charge configuration due to changes in $g$ within each QD modifying $\Delta E_{\mathrm{Z}}$ as QD1 and QD2 are shaped by applied voltages, and the local interface disorder sampled by the electron varies Jock2018 . The charge transition between (3,1,0) and (3,0,1) appears as a diagonal strip of constant $\sim$50singlet return probability indicating a fully dephased qubit. A shift in coherent singlet return on either side of the transition indicates the spin is transferred from one isolated QD to the other (see discussion on Supplementary Fig. 2 for more details). The low singlet return probability and absence of acquired phase bands in (3,1,1) ((3,0,0)) is due to the addition (subtraction) of an electron to the system which rapidly destroys qubit coherence. Altogether, these four delineated regions map out the charge space where spin shuttling operations will be performed and show that the entangled electron spin may be coherently transferred between QD1 and QD2.
@@ II.2 Spin coherence and sources of dephasing
We first consider dephasing in the case of negligible external magnetic field ($B_{0}\approx 0$ mT, with a remanent field of up to a few mT due to trapped flux in the superconducting magnet) such that dephasing is expected to be dominated by fluctuations in the local Overhauser field Petta2005 , where no $S$-$T_{0}$ rotations are observed. We fit the data to a decay profile of the form
@@
where $A$ is the amplitude, $y_{0}$ is the vertical offset, $T_{2}^{*}$ is the inhomogeneous dephasing time, and $\alpha$ is the decay exponent. The resulting dephasing times and fit uncertainties are $T_{2}^{*}(3,1,0)=4.60\pm 0.02\leavevmode\nobreak\ \mu$s, $T_{2}^{*}(3,0,1)=3.86\pm 0.02\leavevmode\nobreak\ \mu$s, and $T_{2}^{*}(\mathrm{transition})=5.12\pm 0.03\leavevmode\nobreak\ \mu$s, and are consistent with previously measured quantities for QDs in 800 ppm 29Si Witzel22012 ; Eng2015 ; Neyens2024 . The variation among $T_{2}^{*}$s suggests local sources of dephasing affect the spin in each QD, including the wavefunction overlap with nearby 29Si and 73Ge. We interpret the longer $T_{2}^{*}(\mathrm{transition})$ as the result of a process like motional narrowing – effectively averaging out magnetic noise acquired from the different nuclear distributions when the spin is hybridized between QD1 and QD2 Mortemousque2021 . This increase in $T_{2}^{*}$ when the second electron interacts with a larger number of nuclear spins when hybridized is consistent with the central limit theorem, which scales as $T_{2}^{*}\sim\sqrt{N_{\mathrm{spins}}}$, where $N_{\mathrm{spins}}$ is the number of spinful nuclei lying within the electron wavefunction Witzel2012 . The resulting decay exponent fits, $\alpha(3,1,0)=1.92\pm 0.02$, $\alpha(3,0,1)=2.16\pm 0.02$, and $\alpha(\mathrm{transition})=2.12\pm 0.02$, resemble Gaussian profiles as expected of inhomogeneous nuclear spin dephasing Petta2005 .
Next, we apply a small external field of $B_{0}=50$ mT to amplify and study the effect of charge noise acting on $\Delta E_{\mathrm{Z}}$. Clear $S$-$T_{0}$ oscillations are observed and fit to a decaying sine function of the form
@@
where $f$ is the $S$-$T_{0}$ rotation frequency. The resulting dephasing times are $T_{2}^{*}(3,1,0)=3.82\pm 0.12\leavevmode\nobreak\ \mu$s, $T_{2}^{*}(3,0,1)=4.09\pm 0.03\leavevmode\nobreak\ \mu$s, and $T_{2}^{*}(\mathrm{transition})=1.08\pm 0.01\leavevmode\nobreak\ \mu$s. We attribute the fast decay observed at the charge transition to the hybridization of the electron wavefunction split between QD1 and QD2 Yoneda2021 ; vanRiggelen-Doelman2024 , which enhances the qubit sensitivity to charge noise. Here, charge noise rapidly alters the extent of the wavefunction overlap within QD1 and QD2, each with its own distinct spin-orbit interactions. This manifests as increased fluctuations in $\Delta E_{\mathrm{Z}}$ experienced by the qubit, which increases with magnetic field, driving faster $S$-$T_{0}$ dephasing. This sensitivity to charge noise may be minimized by tuning the external magnetic field angle Tanttu2019 , or by simply reducing the magnitude. For the remainder of this work concerning shuttling, we maintain a small or negligible field to reduce the susceptibility of the qubit to noise through $\Delta E_{\mathrm{Z}}$.
Additionally, we observe changes to $\alpha$ that imply changes to the noise spectrum. The resulting decay exponent fits are $\alpha(3,1,0)=0.90\pm 0.06$, $\alpha(3,0,1)=2.12\pm 0.06$, and $\alpha(\mathrm{transition})=1.14\pm 0.02$, ranging from exponential to Gaussian forms. Since the noise spectrum and mechanisms dictate the observed decay profile Stano2022 , the various $\alpha$ suggest different localized noise dynamics, in addition to residual nuclear spins, contribute to dephasing in each QD. One possible source known to exhibit non-Gaussian decay behavior is random telegraph noise originating from charge traps, where the observed charge noise is often attributed to two-level systems (TLSs) that are distributed nonuniformly within the gate dielectric Connors2019 ; Struck2020 . This is consistent with our assessment that charge noise dominates the fast $T_{2}^{*}(\mathrm{transition})$, and suggests an increased susceptibility to charge noise in QD1 compared to QD2, since $\alpha(3,1,0)$ is exponential and $\alpha(3,0,1)$ is Gaussian. Overall, we observe qubit inhomogeneous dephasing to be primarily driven by fluctuations in the Overhauser fields generated by residual nuclear spins near 0 mT, and a combination of nuclear spins and charge noise at 50 mT.
Figure 3: Repeated shuttling experiments. a. Pulse sequence diagram for repeated shuttles. A |$\ket{S}$ state is prepared, undergoes shuttling between QD1 and QD2, and then its final spin state is measured. b. Cartoon representation of the exchange interaction between QD0 and QD1,2 used to carry out a $\pi_{z}$ pulse for dynamical decoupling. c. Singlet return probability after $N$ shuttles (different colored markers) are executed over a range of time. A fit of the form of Eq. 3 is overlaid (black curve). d. Singlet return probability for an experiment similar to c., but with the dynamical decoupling sequence shown in the inset applied. A fit of the form of Eq. 5 is overlaid (black curve). e. Singlet return probability for a $\ket{\uparrow\downarrow}$ state after $N$ shuttles executed over time. An exponential decay fit of the form of Eq. 6 is overlaid (black curve). f., g., h. Repeated shuttles carried out at fixed $t_{\mathrm{shuttle}}$ corresponding to the colored vertical dashed lines in c., d., and e., respectively. $B_{0}\approx 0$ mT for the |$\ket{S}$ shuttles (c.,d.,f.,g.) and $B_{0}=10$ mT for the $\ket{\uparrow\downarrow}$ shuttles (e., h.). All error bars shown represent the standard deviation of the mean of the measurement.